# Exploiting Synergistic Cognitive Biases to Bypass Safety in LLMs

### v1

@@ Appendix A Detailed Information
Redteam Model DeepSeek-R1-Distill-Qwen-14B (Guo et al. 2025)
@@
PPO Victim Model Llama-2-7B (Touvron and et al. 2023), Qwen-turbo, GPT-4o-mini, and Deepseek-r1-distill-qwen-7b (Guo et al. 2025)
@@
Training Datasets 50% AdvBench (Zou et al. 2023) and 50% HEx-PHI (Qi et al. 2023)
@@
Evaluation Datasets In-distribution: Remaining 50% of AdvBench + HEX-PHI; Out-of-distribution: HarmBench (Mazeika et al. 2024)
@@
Victim LLM (30) Llama-2-7B, Llama-2-13B, Llama-2-70B, Llama-3.3-70B, Llama-4-maverick-17B, Vicuna-7B, Vicuna-13B, Mistral-7B, Mistral-8$\times$7B), Qwen(Qwen-7B, Qwen-14B, Qwen-turbo, Qwen-plus, Qwen-max, QwQ-32B, GPT-4o-mini, GPT-4o, GPT-4.1, O1-mini, O3-mini, O4-mini, Claude-3-haiku, Gemini-1.5-flash, DeepSeek-v3-241226, DeepSeek-v3-250324, DeepSeek-r1, DeepSeek-r1-distill-qwen-7B, DeepSeek-r1-distill-qwen-32B, DeepSeek-r1-distill-llama-8B, DeepSeek-r1-distill-llama-70B
Best-of-N 3 (Align to HarmBench(Mazeika et al. 2024))
@@
Human Jailbreaks (Shen et al. 2024): Carefully handcrafted prompts that semantically conceal malicious intent to induce policy-violating responses.
GCG (Zou et al. 2023): A white‑box gradient-based attack that iteratively adjusts adversarial suffix tokens to evade alignment constraints.
AutoDAN (Liu et al. 2023): A white‑box evolutionary algorithm that generates readable, concealed adversarial suffixes using a hierarchical genetic approach.
Cognitive Overload (Xu et al. 2023): A black‑box method that overloads the model’s “cognitive” capacity via multilingual confusion, obfuscation, and causal inversion. Proven successful on benchmarks including AdvBench and MasterKey, and effective against models like Llama-2 and ChatGPT.
DeepInception (Li et al. 2023): A black‑box attack inspired by psychology experiments; it constructs nested, hypnotic narratives that deceive the model into self‑hypnosis and repeated jailbreak. It shows high success rates across Llama, Falcon, Vicuna, and GPT.
PAIR (Chao et al. 2023): A black‑box method that iteratively refines prompts over a few shots, efficiently converging on a jailbreak‑capable input.
ArtPrompt (Jiang et al. 2024): A black‑box technique that leverages ASCII‑art encoding to hide malicious payloads within textual puzzles, prompting the model to decode and inadvertently bypass alignment.
PAP (Zeng et al. 2024): A black‑box approach combining prompt taxonomy with persuasive rhetoric, producing coherent and socially engineered adversarial prompts that effectively bypass safety filters.
@@
To quantify the harmfulness of a model’s response, we utilize a GPT-Judge (Qi et al. 2023). This judge [^1] https://github.com/uw-nsl/ArtPrompt/blob/e21a123d4ebb02bb438b23e6edd5ed4b4dadb221/jailbreak/safe˙eval.py#L161 rates each model response on a 1–5 scale, where a score of 5 indicates highly harmful content (e.g., generating dangerous instructions, promoting illegal activities, or producing hate speech), and a score of 1 signals a successful refusal or harmless output (e.g., a statement indicating inability to comply with the request or a benign response). This metric is widely adopted in the field, with representative baselines (Jiang et al. 2024; Chao et al. 2023; Zeng et al. 2024; Li et al. 2023) also employing it to measure the harmfulness of LLM’s responses. A higher harmfulness score indicates a more successful adversarial attack.
@@
The Helpfulness Rate (HPR) (Jiang et al. 2024; Zou et al. 2023; Liu et al. 2023) serves as a crucial indicator of the target model’s robustness against adversarial prompts. It specifically measures the proportion of cases where the target model fails to output rejection responses when presented with an adversarial prompt. This metric directly reflects the model’s susceptibility to generating undesirable content rather than adhering to its safety guidelines. We consider a response to be a ”rejection” if it contains any of the following common refusal phrases [^2] https://github.com/uw-nsl/ArtPrompt/blob/e21a123d4ebb02bb438b23e6edd5ed4b4dadb221/jailbreak/safe˙eval.py#L13:
@@
The Intention Score(Lee et al. 2025) evaluates the degree to which the malicious intent of an original prompt is preserved in its paraphrased or perturbed variant. A judge (typically another LLM or a human expert) conducts a pairwise comparison between the original prompt and the modified prompt. The judge then assigns one of three scores: -1, 0, or 1. This metric is vital for understanding whether an adversarial attack maintains its core malicious objective, even when the surface form of the prompt changes. The detailed system prompt provided to the judge is as Figure 7.
@@
The ASR is our primary metric for evaluating the overall effectiveness of an adversarial attack. It quantifies the percentage of harmful instructions that successfully bypass the LLM’s safety mechanisms. We define ASR based on the Harmfulness Score (HS), which is evaluated by a GPT-Judge (Qi et al. 2023). The HS assesses the degree of harmfulness in the LLM’s output on a scale of 1 to 5, where 5 indicates highly harmful content. Specifically, the ASR is calculated as:
$\textbf{ASR}=\frac{\text{\# of responses with HS = 5}}{\text{\# of total responses}}$
@@ Appendix B Cognitive Bias Paraphraser
This section presents a structured taxonomy of cognitive biases, adapted from the comprehensive work by Dimara et al. (Dimara et al. 2018). This taxonomy categorizes biases based on the primary cognitive TASK they influence (e.g., estimation, decision, recall) and further classifies them by their underlying Flavor (e.g., association, baseline, outcome, self-perspective, inertia). This multi-dimensional classification helps in identifying the root causes of biased thinking and developing targeted interventions to mitigate their effects.
@@
Table 8: Taxonomy of Cognitive Biases (Dimara et al. 2018)

### v2

@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. U24A20335). We thank the shepherd and all the anonymous reviewers for their constructive feedback.
Ethics Statement
@@ Appendix A Detailed Information
Redteam Model DeepSeek-R1-Distill-Qwen-14B (guo2025deepseek)
@@
PPO Victim Model Llama-2-7B (touvron2023llama), Qwen-turbo, GPT-4o-mini, and Deepseek-r1-distill-qwen-7b (guo2025deepseek)
@@
Training Datasets 50% AdvBench (zou2023universal) and 50% HEx-PHI (anonymous2024finetuning)
@@
Evaluation Datasets In-distribution: Remaining 50% of AdvBench + HEX-PHI; Out-of-distribution: HarmBench (mazeika2024harmbench)
@@
[…]
Best-of-N 3 (Align to HarmBench(mazeika2024harmbench))
@@
Human Jailbreaks (shen2024anything): Carefully handcrafted prompts that semantically conceal malicious intent to induce policy-violating responses.
GCG (zou2023universal): A white‑box gradient-based attack that iteratively adjusts adversarial suffix tokens to evade alignment constraints.
AutoDAN (liu2023autodan): A white‑box evolutionary algorithm that generates readable, concealed adversarial suffixes using a hierarchical genetic approach.
Cognitive Overload (xu2023cognitive): A black‑box method that overloads the model’s “cognitive” capacity via multilingual confusion, obfuscation, and causal inversion. Proven successful on benchmarks including AdvBench and MasterKey, and effective against models like Llama-2 and ChatGPT.
DeepInception (li2023deepinception): A black‑box attack inspired by psychology experiments; it constructs nested, hypnotic narratives that deceive the model into self‑hypnosis and repeated jailbreak. It shows high success rates across Llama, Falcon, Vicuna, and GPT.
PAIR (chao2023jailbreaking): A black‑box method that iteratively refines prompts over a few shots, efficiently converging on a jailbreak‑capable input.
ArtPrompt (jiang2024artprompt): A black‑box technique that leverages ASCII‑art encoding to hide malicious payloads within textual puzzles, prompting the model to decode and inadvertently bypass alignment.
PAP (zeng2024johnny): A black‑box approach combining prompt taxonomy with persuasive rhetoric, producing coherent and socially engineered adversarial prompts that effectively bypass safety filters.
@@
To quantify the harmfulness of a model’s response, we utilize a GPT-Judge (qi2023fine). This judge [^1] https://github.com/uw-nsl/ArtPrompt/blob/e21a123d4ebb02bb438b23e6edd5ed4b4dadb221/jailbreak/safe˙eval.py#L161 rates each model response on a 1–5 scale, where a score of 5 indicates highly harmful content (e.g., generating dangerous instructions, promoting illegal activities, or producing hate speech), and a score of 1 signals a successful refusal or harmless output (e.g., a statement indicating inability to comply with the request or a benign response). This metric is widely adopted in the field, with representative baselines (jiang2024artprompt; chao2023jailbreaking; zeng2024johnny; li2023deepinception) also employing it to measure the harmfulness of LLM’s responses. A higher harmfulness score indicates a more successful adversarial attack.
@@
The Helpfulness Rate (HPR) (jiang2024artprompt; zou2023universal; liu2023autodan) serves as a crucial indicator of the target model’s robustness against adversarial prompts. It specifically measures the proportion of cases where the target model fails to output rejection responses when presented with an adversarial prompt. This metric directly reflects the model’s susceptibility to generating undesirable content rather than adhering to its safety guidelines. We consider a response to be a ”rejection” if it contains any of the following common refusal phrases [^2] https://github.com/uw-nsl/ArtPrompt/blob/e21a123d4ebb02bb438b23e6edd5ed4b4dadb221/jailbreak/safe˙eval.py#L13:
@@
The Intention Score(Lee2025xJailbreakRS) evaluates the degree to which the malicious intent of an original prompt is preserved in its paraphrased or perturbed variant. A judge (typically another LLM or a human expert) conducts a pairwise comparison between the original prompt and the modified prompt. The judge then assigns one of three scores: -1, 0, or 1. This metric is vital for understanding whether an adversarial attack maintains its core malicious objective, even when the surface form of the prompt changes. The detailed system prompt provided to the judge is as Figure 6.
@@
The ASR is our primary metric for evaluating the overall effectiveness of an adversarial attack. It quantifies the percentage of harmful instructions that successfully bypass the LLM’s safety mechanisms. We define ASR based on the Harmfulness Score (HS), which is evaluated by a GPT-Judge (qi2023fine). The HS assesses the degree of harmfulness in the LLM’s output on a scale of 1 to 5, where 5 indicates highly harmful content. Specifically, the ASR is calculated as:
[…]
@@ Appendix B Cognitive Bias Paraphraser
This section presents a structured taxonomy of cognitive biases, adapted from the comprehensive work by Dimara et al. (dimara2018task). This taxonomy categorizes biases based on the primary cognitive TASK they influence (e.g., estimation, decision, recall) and further classifies them by their underlying Flavor (e.g., association, baseline, outcome, self-perspective, inertia). This multi-dimensional classification helps in identifying the root causes of biased thinking and developing targeted interventions to mitigate their effects.
@@
Table 9: Taxonomy of Cognitive Biases (dimara2018task)